# Average Certified Radius is a Poor Metric for Randomized Smoothing

### v1

@@ Abstract
Randomized smoothing is a popular approach for providing certified robustness guarantees against adversarial attacks, and has become a very active area of research. Over the past years, the average certified radius (ACR) has emerged as the single most important metric for comparing methods and tracking progress in the field. However, in this work, we show that ACR is an exceptionally poor metric for evaluating robustness guarantees provided by randomized smoothing. We theoretically show not only that a trivial classifier can have arbitrarily large ACR, but also that ACR is much more sensitive to improvements on easy samples than on hard ones. Empirically, we confirm that existing training strategies that improve ACR reduce the model’s robustness on hard samples. Further, we show that by focusing on easy samples, we can effectively replicate the increase in ACR. We develop strategies, including explicitly discarding hard samples, reweighing the dataset with certified radius, and extreme optimization for easy samples, to achieve state-of-the-art ACR, although these strategies ignore robustness for the general data distribution. Overall, our results suggest that ACR has introduced a strong undesired bias to the field, and better metrics are required to holistically evaluate randomized smoothing.
@@ 1 Introduction
Adversarial robustness, namely the ability of a model to resist arbitrary small perturbations to its input, is a critical property for deploying machine learning models in security-sensitive applications. Due to the incompleteness of adversarial attacks which try to construct a perturbation that manipulates the model (Athalye et al., 2018), certified defenses have been proposed to provide robustness guarantees. While deterministic certified defenses (Gowal et al., 2018; Mirman et al., 2018; Shi et al., 2021; Müller et al., 2023; Mao et al., 2023; 2024a; Palma et al., 2023; Balauca et al., 2024) incur no additional cost at inference-time, randomized certified defenses scale better with probabilistic guarantees at the cost of multiplied inference-time complexity. The most popular randomized certified defense is Randomized Smoothing (RS) (Lécuyer et al., 2019; Cohen et al., 2019), which computes the maximum certified radius for every input given the accuracy of a base model on noisy inputs $p_{A}$.
Figure 1: Conceptual illustration of the effect of RS training strategies.
To train better models with larger certified radius under RS, many training strategies have been proposed (Cohen et al., 2019; Salman et al., 2019; Jeong & Shin, 2020; Zhai et al., 2020; Jeong et al., 2023). Average Certified Radius (ACR), defined to be the average of the certified radiuses over each sample in the dataset, has been the main metric to evaluate the effectiveness of these methods. However, in this work, we show that ACR is a poor metric for evaluating the true robustness of a given model under RS. We prove theoretically that ACR of a trivial classifier could be arbitrarily large given enough certification budget, and then show empirically that state-of-the-art (SOTA) RS training strategies reduce the accuracy on hard inputs to increase the ACR. Further, we demonstrate that through explicitly reweighing the training data to focus only on easy inputs, the simplest Gaussian training can be gradually amplified to achieve a SOTA ACR, questioning the development of RS training strategies.
@@
  • Based on these novel insights, we develop strategies to amplify Gaussian training to achieve a SOTA ACR by reweighing the training data to focus only on easy inputs. Specifically, we discard hard inputs during training, reweigh the dataset with their contribution to ACR, and push $p_{A}$ extremely close to 1 for easy inputs via adversarial noise selection. With these simple modifications to Gaussian training, which do not optimize robustness for the general data distribution, but only ACR, we achieve a new SOTA in ACR (Section 5 and Section 6).
Overall, our work proves the need for new metrics for RS. In particular, we suggest to use certified accuracy at various radii as a more informative metric and encourage the community to re-evaluate existing RS training more uniformly (Section 7). We hope this work can inspire future research in this direction.
@@ 3 Background
Randomized Smoothing constructs a smooth classifier $\hat{f}(x)$ given a base classifier $f$, defined as follows: $\hat{f}(x):=\operatorname*{arg\,max}_{c\in{\mathcal{Y}}}\mathbb{P}_{\delta\sim{\mathcal{N}}(0,\sigma^{2}I)}(f(x+\delta)=c)$. Intuitively, the smooth classifier assigns the label with maximum probability in the neighborhood of the input. With this formulation, Cohen et al. (2019) proves that $\hat{f}$ is adversarially robust within $B_{R(x,p_{A})}(x)$ when $p_{A}\geq 0.5$, where $R(x,p_{A}):=\sigma\Phi^{-1}(p_{A})$, $\Phi$ is the cumulative distribution function of ${\mathcal{N}}(0,1)$ and $p_{A}:=\max_{c\in{\mathcal{Y}}}\mathbb{P}_{\delta\sim{\mathcal{N}}(0,\sigma^{2}I)}(f(x+\delta)=c)$ is the probability of the most likely class. Average Certified Radius (ACR) is defined as the average of $R(x,p_{A})I(\hat{f}(x)=y)$ over the dataset. In practice, $p_{A}$ cannot be computed exactly, and an estimation $\hat{p}_{A}$ such that $\mathbb{P}(p_{A}\geq\hat{p}_{A})\geq 1-\alpha$ is substituted, where $\alpha$ is the confidence threshold and $\hat{p}_{A}$ is computed based on $N$ trials for the event $I(f(x+\delta)=c)$. We call $N$ the certification budget, which is the number of queries to the base classifier $f$ to estimate $p_{A}$. Since RS certifies robustness based on the accuracy of the base model on samples perturbed by Gaussian noise, Gaussian training, which augments the train data with Gaussian noise, is the most common method to train the base model for RS. Specifically, it optimizes
@@
where $\delta_{i}$ are sampled from ${\mathcal{N}}(0,\sigma^{2}I)$ uniformly at random, $m$ is the number of samples and $L$ is the cross entropy loss.
@@ 4.1 Trivial Classifier with Infinite ACR
Proof.
Theorem 1 shows that ACR can be arbitrarily large for a trivial classifier with large enough certification budget. This implies that ACR is not reliable for evaluating a model under RS, as a trivial classifier can achieve infinite ACR with minimal robustness on at least half of classes. In practice, the budget certification $N$ is usually limited, and thus $R(x,p_{A})$ is bounded by a constant for every $x$ and $p_{A}$. In this case, the ACR of a trivial classifier is also bounded. However, in Section 4.2 below, we will show that this is still problematic, as ACR is much more sensitive to improvements on easy samples than on hard samples.
@@ 4.2 ACR Strongly Prefers Easy Samples
We now discuss the effect of ACR with a limited budget. We follow the standard certification setting in the literature, setting $N=10^{5}$ and $\alpha=0.001$. With this budget, the maximum certified radius for one input is $R(x,p_{A}=1)=\sigma\Phi^{-1}(\alpha^{1/N})\approx 3.8\sigma$. However, we will show that $\frac{\partial R(x,p_{A})}{\partial p_{A}}$ grows extremely fast, exceeding $1000\sigma$ when $p_{A}\rightarrow 1$ and close to 00 when $p_{A}\rightarrow 0.5$.
Without loss of generality, we set $\sigma=1$ and denote $R(x,p_{A})$ as $r$. Figure 2 shows $r$ and $\frac{\partial r}{\partial p_{A}}$ against $p_{A}$. While $r$ remains bounded by a constant $3.8$, $\frac{\partial r}{\partial p_{A}}$ grows extremely fast when $p_{A}$ approaches 1. As a result, increasing $p_{A}$ from $0.99$ to $0.999$ improves $r$ from $2.3$ to $3.0$, matching the improvement achieved by increasing $p_{A}$ from 00 to $0.76$. Therefore, to achieve maximum ACR, it is much more efficient for the training algorithm to focus on improving $p_{A}$ on easy samples where $p_{A}$ is close to 1. Further, when $p_{A}<0.5$, the data point will not contribute to ACR at all, thus optimizing ACR will not increase $p_{A}$ with a local optimization algorithm like gradient descent. Therefore, it is natural for RS training to disregard inputs with $p_{A}<0.5$ as their ultimate goal is to improve ACR.
4.3 RS Training Trades off Hard Samples for ACR
We have shown that ACR strongly prefers easy samples in Section 4.2. However, since ACR is not differentiable with respect to the model parameters because it is based on counting, RS training strategies usually do not directly apply ACR as the training loss. Instead, they optimize various surrogate objectives, and finally evaluate the model with ACR. Thus, it is unclear whether and to what extent the design of training algorithms is affected by the ACR metric. We now empirically quantify the effect, confirming the theoretical analysis. Specifically, we show that SOTA training strategies reduce $p_{A}$ of hard samples and put more weight (measured by gradient norm) on easy samples compared to Gaussian training.
Figure 3 shows the empirical cumulative distribution of $p_{A}$ for models trained with SOTA algorithms and Gaussian training. Clearly, for various $\sigma$, SOTA algorithms have higher density than Gaussian training at $p_{A}$ close to zero and one. While they gain more ACR due to the improvement on easy samples, hard samples are consistently underrepresented in the final model compared to Gaussian training. As a result, Gaussian training has higher $\mathbb{P}(p_{A}\geq 0.5)$ (clean accuracy), and SOTA algorithms exceed Gaussian training only when $p_{A}$ passes a certain threshold, i.e., when the certified radius is relatively large. This is problematic in practice, indicating that ACR does not properly measure the model’s ability. For example, a face recognition model could have a high ACR but consistently refuse to learn some difficult faces, which is not acceptable in real-world applications.
@@ 4.3 RS Training Trades off Hard Samples for ACR
Figure 3: The empirical cumulative distribution of $p_{A}$ on CIFAR-10 for models trained and certified with various $\sigma$ with different training algorithms.
To further quantify the relative weight between easy and hard samples indicated by each training algorithm, we measure the average gradient $l_{2}$ norm of easy and hard samples for models trained with different algorithms and $\sigma=0.5$, as a proxy for the sample weight. Intuitively, samples with larger gradients contribute more to training and thus are more important for the final model. As shown in Table 1, Gaussian training puts less weight on easy samples than hard samples, which is natural as easy samples have smaller loss values. However, SOTA algorithms put more weight on easy samples compared to Gaussian training, e.g., the relative weight between easy and hard samples is 4.28 for CAT-RS (Jeong et al., 2023), while for Gaussian training it is 0.45. This confirms that SOTA algorithms indeed prioritize easy samples over hard samples, consistent with the theoretical analysis in Section 4.2.
@@
Table 1: The average gradient $l_{2}$ norm of easy ($p_{A}>0.5$) and hard ($p_{A}<0.5$) samples for models trained with different algorithms and $\sigma=0.5$, along with their relative magnitude (easy / hard). The corresponding ACR is also shown.
5 Amplifying Easy Data Greatly Improves ACR
In Section 4, we concluded that ACR strongly prefers easy samples, and RS training trades off hard samples for ACR. This raises the question of whether explicitly focusing on easy data during training can effectively replicate the increase in ACR. In this section, we propose three modifications to the simplest Gaussian training to achieve this goal.
@@ 5.1 Discard Hard Data During Training
5.2 Data Reweighing with Certified Radius
@@
ACR relates non-linearly to $p_{A}$, and the growth of the certified radius is much faster for easy samples with high $p_{A}$ (Figure 2). We account for this by reweighing the data points based on their certified radius. Specifically, we use the approximate (normalized) certified radius as the weight of the probability for every data point being sampled. We formulate the weight $w$ of every data point $x$ as:
@@
where $C$ is the count of correctly classified noisy samples and $p_{\rm min}$ is the reference probability threshold. Note that the estimation of $\hat{p}_{A}$ aligns with the certification. We normalize the weight to be at least 1 since the original radius is zero when $p_{A}$ is relatively low. To minimize computational overhead, we evaluate $\hat{p}_{A}$ every 10 epoch with $N=16$ and $\alpha=0.1$ throughout the paper. We set $p_{\rm min}=0.75$ because this is the minimum probability that has positive certified radius under this setting. The sampling weight curve is visualized in Figure 4. After the reweighing, ACR has an approximately linear relationship to $p_{A}$ when $w>1$, and easy samples are sampled more frequently than hard samples so that Gaussian training also improves $p_{A}$ further for easy samples.

### v2

@@ Abstract
Randomized smoothing is a popular approach for providing certified robustness guarantees against adversarial attacks, and has become an active area of research. Over the past years, the average certified radius (ACR) has emerged as the most important metric for comparing methods and tracking progress in the field. However, in this work, for the first time we show that ACR is a poor metric for evaluating robustness guarantees provided by randomized smoothing. We theoretically prove not only that a trivial classifier can have arbitrarily large ACR, but also that ACR is much more sensitive to improvements on easy samples than on hard ones. Empirically, we confirm that existing training strategies, though improving ACR with different approaches, reduce the model’s robustness on hard samples consistently. To strengthen our conclusion, we propose strategies, including explicitly discarding hard samples, reweighting the dataset with approximate certified radius, and extreme optimization for easy samples, to achieve state-of-the-art ACR, without training for robustness on the full data distribution. Overall, our results suggest that ACR has introduced a strong undesired bias to the field, and its application should be discontinued when evaluating randomized smoothing.
Machine Learning, ICML
[…]
return
@@ 1 Introduction
Adversarial robustness, namely the ability of a model to resist arbitrary small perturbations to its input, is a critical property for deploying machine learning models in security-sensitive applications. Due to the incompleteness of adversarial attacks which try to construct a perturbation that manipulates the model (Athalye et al., 2018), certified defenses have been proposed to provide robustness guarantees. While deterministic certified defenses (Gowal et al., 2018; Mirman et al., 2018; Shi et al., 2021; Müller et al., 2023; Mao et al., 2023, 2024a; Palma et al., 2023; Balauca et al., 2024) incur no additional cost at inference-time, randomized certified defenses scale better with probabilistic guarantees at the cost of multiplied inference-time complexity. The most popular randomized certified defense is Randomized Smoothing (RS) (Lécuyer et al., 2019; Cohen et al., 2019), which computes the maximum certified radius for every input given the accuracy of a base model on noisy inputs $p_{A}$.
Figure 1: Conceptual illustration of the effect of RS training.
To train better models with larger certified radius under RS, many training strategies have been proposed (Cohen et al., 2019; Salman et al., 2019; Jeong & Shin, 2020; Zhai et al., 2020; Jeong et al., 2023). Average Certified Radius (ACR), defined to be the average of the certified radiuses over each sample in the dataset, has been the main metric to evaluate the effectiveness of these methods. Although some studies refrained from using ACR in their evaluation unintentionally (Carlini et al., 2022; Xiao et al., 2022), the inherent limitation of ACR remains unknown. As a result, the community remains largely unaware of these issues, resulting in subsequent research relying on ACR continuously (Zhang et al., 2023; Jeong & Shin, 2024). However, in this work, we show that ACR is a poor metric for evaluating the true robustness of a given model under RS.
@@
  • Based on these novel insights, we develop strategies to amplify Gaussian training to achieve the SOTA ACR by reweighting the training data to focus only on easy inputs. Specifically, we discard hard inputs during training, reweight the dataset with their contribution to ACR, and push $p_{A}$ extremely close to 1 for easy inputs via adversarial noise selection. With these simple modifications to Gaussian training, we achieve the new SOTA in ACR without targeting robustness for the general case (Section 5 and Section 6).
Overall, our work proves that the application of ACR should be discontinued in the evaluation of randomized smoothing. In particular, we suggest using the best certified accuracy at various radii and the empirical distribution of $p_{A}$ as more informative metrics, and encourage the community to evaluate RS training more uniformly (Section 7). We hope this work will motivate the community to permanently abandon ACR and inspire future research in proper evaluation of randomized smoothing.
@@ 3 Background
Randomized Smoothing constructs a smooth classifier $\hat{f}(x)$ given a base classifier $f$, defined as follows: $\hat{f}(x):=\operatorname*{arg\,max}_{c\in{\mathcal{Y}}}\mathbb{P}_{\delta\sim{\mathcal{N}}(0,\sigma^{2}I)}(f(x+\delta)=c)$. Intuitively, the smooth classifier assigns the label with maximum probability in the neighborhood of the input. With this formulation, Cohen et al. (2019) proves that $\hat{f}$ is adversarially robust within $B_{R(x,p_{A})}(x)$ when $p_{A}\geq 0.5$, where $R(x,p_{A}):=\sigma\Phi^{-1}(p_{A})$, $\Phi$ is the cumulative distribution function of ${\mathcal{N}}(0,1)$ and $p_{A}:=\max_{c\in{\mathcal{Y}}}\mathbb{P}_{\delta\sim{\mathcal{N}}(0,\sigma^{2}I)}(f(x+\delta)=c)$ is the probability of the most likely class. Average Certified Radius (ACR) is defined as the average of $R(x,p_{A})I(\hat{f}(x)=y)$ over the dataset. In practice, $p_{A}$ cannot be computed exactly, and an estimation $\hat{p}_{A}$ such that $\mathbb{P}(p_{A}\geq\hat{p}_{A})\geq 1-\alpha$ is substituted, where $\alpha$ is the confidence threshold and $\hat{p}_{A}$ is computed based on $N$ trials for the event $I(f(x+\delta)=c)$. We call $N$ the certification budget, which is the number of queries to the base classifier $f$ to estimate $p_{A}$. Since RS certifies robustness based on the accuracy of the base model on samples perturbed by Gaussian noise, Gaussian training, which augments the train data with Gaussian noise, is the most intuitive method to train the base model for RS. Specifically, it optimizes
@@
where $\delta_{i}$ are sampled from ${\mathcal{N}}(0,\sigma^{2}I)$ uniformly at random, $m$ is the number of samples and $L$ is the cross entropy loss. Due to space constraints, we omit the details of other training strategies for RS, and refer interested readers to Cohen et al. (2019); Salman et al. (2019); Jeong & Shin (2020); Zhai et al. (2020); Jeong et al. (2023).
@@ 4.1 Trivial Classifier with Infinite ACR
The proof of Theorem 1 is deferred to Section A.1. Despite an intuitive proof, Theorem 1 reveals an important fact: ACR can be arbitrarily large for a classifier with large enough certification budget. Further, a trivial classifier can achieve infinite ACR with minimal robustness on at least half of classes. Therefore, the ACR metric presented in the literature is heavily associated with the certification budget; when the certification budget changes, the order of ACR between models may change as well, since inputs with lower $p_{A}$ require a larger budget to certify. For example, given a perfectly balanced binary dataset and $N=50$, $\alpha=0.001$, $\sigma=1$, a trivial classifier can achieve ACR equals $0.565$, while another classifier with $p_{A}=0.9$ on every input can only achieve ACR equals $0.544$, making the trivial classifier the better one. However, when $N$ is increased to $100$, the latter achieves ACR equals $0.756$, surpassing the trivial classifier which has ACR equals $0.750$. Interestingly, when we further increase $N$ to $200$, the trivial classifier achieves ACR equals $0.913$, surpassing the latter which has ACR equals $0.909$, thus the order is reversed again. This toy example showcases the strong dependence of the evaluation results, especially the comparative order of different models, on the certification budget. This is problematic in practice, as the budget usually differs depending on the task and the model, resulting in uncertainty when choosing training strategies evaluated based on ACR.
@@ 4.2 ACR Strongly Prefers Easy Samples
We now discuss the effect of ACR with a constant budget and fixed $\alpha$. We follow the standard certification setting in the literature, setting $N=10^{5}$ and $\alpha=0.001$. With this budget, the maximum certified radius for one input is $R(x,p_{A}=1)=\sigma\Phi^{-1}(\alpha^{1/N})\approx 3.8\sigma$. However, we will show that $\frac{\partial R(x,p_{A})}{\partial p_{A}}$ grows extremely fast, exceeding $1000\sigma$ when $p_{A}\rightarrow 1$ and close to 00 when $p_{A}\rightarrow 0.5$.
Without loss of generality, we set $\sigma=1$ and denote $R(x,p_{A})$ as $r$. Figure 2 shows $r$ and $\frac{\partial r}{\partial p_{A}}$ against $p_{A}$. While $r$ remains bounded by a constant $3.8$, $\frac{\partial r}{\partial p_{A}}$ grows extremely fast when $p_{A}$ approaches 1. As a result, increasing $p_{A}$ from $0.99$ to $0.999$ improves $r$ from $2.3$ to $3.0$, matching the improvement achieved by increasing $p_{A}$ from 00 to $0.76$. Therefore, to achieve maximum ACR, it is much more efficient for the training algorithm to focus on improving $p_{A}$ on easy samples where $p_{A}$ is close to 1. Further, when $p_{A}<0.5$, the data point will not contribute to ACR at all, thus optimizing ACR will not increase $p_{A}$ with a local optimization algorithm like gradient descent. Therefore, it is natural for RS training to disregard inputs with $p_{A}<0.5$, as their ultimate goal is to improve ACR.
We use the same toy example as in Section 4.1 to further demonstrate this problem. If we further increase $N$ beyond $200$, we can find, surprisingly, that the classifier with $p_{A}=0.9$ for every input always has a smaller ACR than the trivial classifier. This means that achieving $p_{A}=1$ on only half of the dataset and getting $p_{A}=0$ for the rest is “more robust” than achieving $p_{A}=0.9$ on every input when evaluated with ACR, which is misaligned with our intuition. This toy example highlights again the strong preference of ACR for improving easy samples, leading to undesired biases.
@@
Figure 3: The empirical cumulative distribution (ECDF) of $p_{A}$ on CIFAR-10 for models trained and certified with various $\sigma$ with different training algorithms.
4.3 Selection Bias in Existing RS Training Algorithms
We have shown that ACR strongly prefers easy samples in Section 4.2. However, since ACR is not differentiable with respect to the model parameters because it is based on counting, RS training strategies usually do not directly apply ACR as the training loss. Instead, they optimize various surrogate objectives, and finally evaluate the model with ACR. Thus, it is unclear whether and to what extent the design of training algorithms is affected by the ACR metric. We now empirically quantify the effect, confirming that the theoretical weakness of ACR has introduced strong selection biases in existing RS training algorithms. Specifically, we show that SOTA training strategies reduce $p_{A}$ of hard samples and put more weight (measured by gradient norm) on easy samples compared to Gaussian training.
Figure 3 shows the empirical cumulative distribution of $p_{A}$ for models trained with SOTA algorithms and Gaussian training. Clearly, for various $\sigma$, SOTA algorithms have higher density than Gaussian training at $p_{A}$ close to zero and one. While they gain more ACR due to the improvement on easy samples, hard samples are consistently underrepresented in the final model compared to Gaussian training. As a result, Gaussian training has higher $\mathbb{P}(p_{A}\geq 0.5)$ (clean accuracy), and SOTA algorithms exceed Gaussian training only when $p_{A}$ passes a certain threshold, i.e., when the certified radius is relatively large. This is problematic in practice, indicating that ACR does not properly measure the model’s ability even with a fixed certification setting. For example, a face recognition model could have a high ACR but consistently refuse to learn some difficult faces, which is not acceptable in real-world applications.
To further quantify the relative weight between easy and hard samples indicated by each training algorithm, we measure the average gradient $l_{2}$ norm of easy and hard samples for models trained with different algorithms and $\sigma=0.5$, as a proxy for the sample weight. Intuitively, samples with larger gradients contribute more to training and thus are more important for the final model. As shown in Table 1, Gaussian training puts less weight on easy samples than hard samples, which is natural as easy samples have smaller loss values. However, SOTA algorithms put more weight on easy samples compared to Gaussian training, e.g., the relative weight between easy and hard samples is 4.28 for the best ACR baseline, CAT-RS (Jeong et al., 2023), while for Gaussian training it is 0.45. This confirms that SOTA algorithms indeed prioritize easy samples over hard samples, consistent with the theoretical analysis in Section 4.2.
@@ 4.3 Selection Bias in Existing RS Training Algorithms
Table 1: The average gradient $l_{2}$ norm of easy ($p_{A}>0.5$) and hard ($p_{A}<0.5$) samples for models trained with different algorithms and $\sigma=0.5$, along with their relative magnitude (easy / hard). The corresponding ACR is provided for reference.
5 Replicating the Progress in ACR
In Section 4, we find that ACR strongly prefers easy samples, and this bias has been introduced to the design of RS training strategies. This makes one wonder the question of whether explicitly focusing on easy data during training can effectively replicate the increase in ACR gained by years spent developing RS training strategies. In this section, we propose three intuitive modifications to the simplest Gaussian training to achieve this goal.
@@ 5.1 Discard Hard Data During Training
5.2 Data Reweighting with Certified Radius
@@
ACR relates non-linearly to $p_{A}$, and the growth of the certified radius is much faster for easy samples. We account for this by reweighting the data points based on their certified radius. Specifically, we use the approximate (normalized) certified radius as the weight of the probability for every data point being sampled. We formulate the weight $w$ of every data point $x$ as:
@@
where $C$ is the count of correctly classified noisy samples and $p_{\rm min}$ is the reference probability threshold. Note that the estimation of $\hat{p}_{A}$ aligns with the certification. We clip the weight at 1 to avoid zero weights. To minimize computational overhead, we evaluate $\hat{p}_{A}$ every 10 epoch with $N=16$ and $\alpha=0.1$ throughout the paper. We set $p_{\rm min}=0.75$ because this is the minimum probability that has positive certified radius under this setting. The sampling weight curve is visualized in Figure 4. After the reweighting, ACR has an approximately linear relationship to $p_{A}$ when $w>1$, and easy samples are sampled more frequently than hard samples so that Gaussian training also improves $p_{A}$ further for easy samples.